# Towards Trapped-Ion Thermometry Using Cavity-Based EIT

### 09 Mar 2026 v3

@@ VII Thermometry of a multi-ion cavity-QED system with a large excited-state decay rate
Figure 7: Cavity-EIT linewidth variation with the number of ions for the case of high decay rate. The simulation parameters are shown in the inset.
A large decay rate induces decoherence in the system, which washes out the cavity-EIT effect, and only the bare cavity signal remains visible at the output. Several groups worldwide have demonstrated the cavity-EIT effect using ion/atom species with a large decay rate from the excited state [44, 19]. The experimental results from these systems show that for a single ion the cavity-EIT signal is not clearly visible; however, increasing the number of atoms leads to the appearance of cavity-EIT peaks. This occurs because, although the ion-cavity system is not configured in the strong coupling regime for a single ion/atom (as the decay rates exceed the cavity coupling factor), the effective coupling strength increases with multiple ions coupled to the intracavity field. Consequently, the collective ion-cavity coupling can overcome the large decay rates of the excited states, as discussed in section VI. More discussion about their system and an excellent agreement of our simulation results with the analytical expression and the reported experimental data can be found in Appendix A and B.
@@
To assess the applicability of the thermometry method in the presence of strong excited-state decay, we analyze the temperature dependence of the cavity-EIT linewidth for different numbers of ions as shown in Figure 8. A distinct and monotonic dependence of the normalized linewidth on the temperature is preserved even for large decay rates ($\gamma_{eu}=20\kappa$, $\gamma_{eg}=7\kappa$). Although the decoherence caused by spontaneous emission, which also partially broadens the EIT feature, collective enhancement ($g_{eff}\propto\sqrt{N}$) strengthens the effective light-matter interaction to compensate for decoherence. These results demonstrate that the suggested cavity-EIT thermometry method is robust even in regimes of significant radiative decay.
VIII Roadmap towards experimental realization
@@
One may also consider using a dipole allowed transition (where typical linewidths are in the range of $\approx 5-20MHz$) for the trapped ion strongly coupled to the cavity field, one should consider operating the ion trap in a very tight confinement (very deep trap potential) of the ion with secular frequency greater than the transition linewidth ($\Gamma$). Such a mode of ion trap operation has been experimentally validated [45], where an 9Be+ ion was trapped with a very high RF voltage (600 V) and frequency (240 MHz), reaching secular frequencies in the range of $\sim 50~MHz$ and the sideband transitions were addressed in a dipole-allowed transition.
@@ Appendix B Comparison of simulation results with analytical expressions and experimental data
With these analytical expression and the parameters listed in [41] we compute the transmission spectrum and compared it with the outcomes of our numerical model. Figure 10 shows the normalized cavity emission as a function of probe detuning $(\Delta/\kappa)$ for $N_{ions}=17$. As observed, the simulation accurately reproduces the key spectral features, including the central transparency peak and two sidebands.
@@
A two-dimensional map of the cavity-EIT linewidth represented as the normalized full width at half maxima (FWHM) ratio as a function of the normalized control field Rabi frequency $\Omega_{c}$ and the normalized atom-cavity coupling strength $g$ for various mean thermal phonon numbers $n_{th}$ is shown in Figure 12 For a weakly excited motional state ($n_{th}=0.5$) (see 12(a)), the linewidth ratio remains close to unity over a large region of the parameter space. This suggests that thermal motion only slightly alters EIT coherence. As the value of thermal occupation increases $n_{th}=1$ (see 12(b)), a clear reduction of the FWHM ratio occurs, particularly in the regime of strong ion-cavity coupling and weak control fields. This reduction signals the onset of thermally induced decoherence. A higher phonon number ($n_{th}=5,10$) (see 12(c), 12(d)) shows a more noticeable trend in which the parameter space is dominated by extended regions of reduced linewidth ratio. In this regime, thermal motion significantly perturbs the ion-cavity interaction, leading to enhanced dephasing and a degradation of the EIT interference.
@@ Appendix D Verification of different sideband transition -related phenomenon
Figure 14: Rabi oscillation in blue sideband transition for $n=0\rightarrow 1$. The red dots represent the data from the NIST experiment [9], and the solid blue curve represents the simulated result obtained using the same parameters as those employed in the experiment. The parameters are, $\eta=0.202$, $\Omega=2\pi\times 500$ kHz, $\gamma=20$ kHz.
The dephasing and frequency in the Rabi oscillation of the blue sideband transition depend on the phonon occupation. There are many experimental verifications for this phenomenon [7, 9]. We have verified this argument, and it matches the experimental data reported in the NIST experiment (see Figure 14) [9].

### 26 Mar 2026 v4

@@ VII Thermometry of a multi-ion cavity-QED system with a large excited-state decay rate
Figure 7: Variation of cavity-EIT linewidth with the number of ions with large decay rate of ion excited state. The simulation parameters are shown in the inset.
A large decay rate induces decoherence in the system, leading to a negligible cavity-EIT effect, and only the bare cavity signal remains visible at the output. Several groups worldwide have demonstrated the cavity-EIT effect using ion/atom species with a large decay rate from the excited state [52, 23]. The experimental results from these systems show that for a single ion the cavity-EIT signal is not clearly visible; however, increasing the number of atoms/ions coupled to the cavity mode leads to the re-appearance of cavity-EIT spectrum. This occurs because, although the ion-cavity system is not configured in the strong coupling regime for a single ion/atom (as the decay rates far exceed the single atom cavity coupling factor), the effective coupling strength increases with multiple ions coupled to the intracavity field. Consequently, the collective ion-cavity coupling can overcome the large decay rates of the excited states, as discussed in section VI. More discussion about their system and an excellent agreement of our simulation results with the analytical expression and the reported experimental data can be found in Appendix A and B.
@@
To assess the applicability of the thermometry method in the presence of an excited state with a relatively large decay rate (>> 7-8 MHz), we analyze the temperature dependence of the cavity-EIT linewidth for different numbers of ions as shown in Figure 8. A distinct and monotonic dependence of the normalized linewidth on the temperature is preserved even for large decay rates ($\gamma_{eu}=20\kappa$, $\gamma_{eg}=7\kappa$). Although the decoherence caused by spontaneous emission, which also partially broadens the EIT feature, collective enhancement ($g_{eff}\propto\sqrt{N}$) strengthens the effective light-matter interaction to compensate for decoherence. These results demonstrate that the suggested cavity-EIT thermometry method is robust even in regimes of significant radiative decay.
VIII Projected sensitivity of cavity-induced EIT thermometry
As discussed in the previous two sections, our method is applicable to both the weak-coupling regime and to systems with large excited-state decay rates. An important question that might arise pertains to how accurately our proposed technique can sense variations in temperature and what is the potential sensitivity corresponding to the average phonon occupation. Addressing this question provides insight into the practical applicability of the method across different parameter regimes. Furthermore, a sensitivity analysis helps to quantify the precision of the thermometry technique and highlights its strengths as well as its limitations.
From the observations in Figure 6 and Figure 8, it is evident that the variation of the linewidth is not uniform across the entire temperature range. Consequently, the sensitivity of our method depends on the specific temperature regime under consideration. To quantify this behavior, we define the temperature sensitivity as the change in temperature corresponding to a unit change in the linewidth, which can be expressed as,
$S_{T}=\frac{\Delta(\Delta\nu)}{\Delta T}$ (12)
where $\Delta\nu$ represents the cavity-EIT linewidth. Similarly the sensitivity for $\bar{n}$ can be written as,
$S_{\bar{n}}=\frac{\Delta(\Delta\nu)}{\Delta\bar{n}}$ (13)
The minimum achievable temperature of the ion, as well as the separation between the motional energy levels, is inherently determined by the secular frequency of the ion trap. This follows from the fact that the effective temperature corresponding to the ground-state energy of the phonon ladder scales with the secular frequency as $T\sim\frac{\hbar\omega_{sec}}{2k_{B}}$. Consequently, a higher secular frequency leads to a larger energy spacing between the motional levels and therefore to a higher equivalent temperature scale associated with these levels. The minimum detectabtle temperature change on the other hand depends on the minimum resolvable linewidth change dictated by the mirror reflectivity.
At higher temperatures, the sensitivity appears to be relatively low, whereas it increases at lower temperatures. To estimate the achievable sensitivity, we assume that the minimum detectable change in the cavity-EIT linewidth is $10~\text{kHz}$.
Figure 9: Temperature vs cavity-EIT linewidth for systems operating in the strong coupling regime for a single ion. The parameters are chosen from [51]. Single ion cavity coupling g=12.3MHzg=12.3~MHz, cavity decay rate, $\kappa=4.1~MHz$, excited state decay, $\gamma=12.5~MHz$. The control field Rabi frequency ($\Omega_{c}$) is taken as 3MHz3~MHz. The secular frequency is taken to be 15MHz15~MHz. The inset plot shows the variation in temperature and average phonon number over a broader range of cavity-EIT linewidth
We have calculated the sensitivity for both the strong and weak-coupling regimes in the presence of a large excited-state decay rate, which represents a realistic and widely encountered experimental scenario. The parameters used in our simulations are inspired by those reported in the experiments of [51] and [38], with the exception of the secular frequency of the ion. In our case, the secular frequency must exceed the decay rates of the excited state in order to satisfy the resolved sideband condition. Accordingly, the secular frequency is chosen to be approximately $15~\text{MHz}$.
Figure 9 represents the single-atom strong-coupling case. Typically, such conditions prevail in fiber cavity-based QED systems. From the slope of the cavity-EIT linewidth versus temperature and $\bar{n}$ plot, we get the value of $14~\mu K/kHz$ and $0.01/kHz$, respectively. But determining $1kHz$ resolution in linewidth can be experimentally challenging, so if we take a minimum of 10kHz10~kHz change in linewidth, the minimum detectable temperature change will be $140~\mu K$ and the minimum detectable deviation in $\bar{n}$ will be $0.1$. As discussed in section VII, the necessary condition for this thermometry method is to be in the resolved sideband regime ($\gamma\ll\omega_{sec}$). For this reason, the secular frequency is taken as 15MHz15~MHz, which is higher than the excited state decay rate of 12.5MHz12.5~MHz.
Figure 10: Temperature vs cavity-EIT linewidth for systems operating in the weak coupling regime for a single ion. The inset plot is for a higher range of temperature, average phonon number and cavity-EIT linewidth. The parameters are chosen the same as the system given in [38]. The single ion cavity coupling factor g=0.54MHzg=0.54~MHz. The cavity decay rate is $\kappa=2.2~MHz$, and the excited state decay rate is $\gamma=12.5~MHz$. The Rabi frequency for control field ($\Omega_{c}$) is 3MHz3~MHz. The secular frequency is considered as 15MHz15~MHz to maintain the resolved sideband regime.
Figure 10 illustrates the sensitivity for the multi-atom low-coupling regime. About $500$ atoms are considered to be coupled to the cavity, resulting in effective strong coupling. The sensitivity in this case is about $120~\mu K$ per 10kHz10~kHz change in linewidth.
The sensitivities in both cases are nearly identical, as both are in an effective strong coupling regime with a similar effective coupling factor. A similar behaviour is also observed when considering the average phonon occupation number ($\bar{n}$).
In our case, as represented in Figure 9 and Figure 10, the assumed secular frequency is ($\sim 15~\mathrm{MHz}$), which is required to satisfy the strict conditions discussed earlier. This secular frequency value is significantly higher than that used in conventional trapped-ion experiments (typically $\sim 1~\mathrm{MHz}$) since the analysis is performed for calcium (40Ca+) ion. As a result, the temperature change corresponding to a given variation in the EIT linewidth, as well as in $\bar{n}$, becomes larger. This reduces the temperature sensitivity of the method, although the sensitivity in terms of the average phonon occupation number $\bar{n}$ remains unaffected.
IX Roadmap towards experimental realization
@@
One may also consider using a dipole allowed transition (where typical linewidths are in the range of $\approx 5-20MHz$) for the trapped ion strongly coupled to the cavity field, one should consider operating the ion trap in a very tight confinement (very deep trap potential) of the ion with a secular frequency greater than the transition linewidth ($\Gamma$). Such a mode of ion trap operation has been experimentally validated [53], where an 9Be+ ion was trapped with a very high RF voltage (600 V) and frequency (240 MHz), reaching secular frequencies in the range of $\sim 50~MHz$ and the sideband transitions were addressed in a dipole-allowed transition.
@@ Appendix B Comparison of simulation results with analytical expressions and experimental data
With these analytical expression and the parameters listed in [47] we compute the transmission spectrum and compared it with the outcomes of our numerical model. Figure 12 highlighted the normalized cavity emission as a function of probe detuning $(\Delta/\kappa)$ having $N_{ions}=17$. As observed, the simulation accurately reproduces the key spectral features, including the central transparency peak and two sidebands.
@@
A two-dimensional map of the cavity-EIT linewidth represented as the normalized full width at half maxima (FWHM) ratio as a function of the normalized control field Rabi frequency $\Omega_{c}$ and the normalized atom-cavity coupling strength $g$ for various mean thermal phonon numbers $n_{th}$ is shown in Figure 14 For a weakly excited motional state ($n_{th}=0.5$) (see 14(a)), the linewidth ratio remains close to unity over a large region of the parameter space. This suggests that thermal motion only slightly alters EIT coherence. As the value of thermal occupation increases $n_{th}=1$ (see 14(b)), a clear reduction of the FWHM ratio occurs, particularly in the regime of strong ion-cavity coupling and weak control fields. This reduction signals the onset of thermally induced decoherence. A higher phonon number ($n_{th}=5,10$) (see 14(c), 14(d)) shows a more noticeable trend in which the parameter space is dominated by extended regions of reduced linewidth ratio. In this regime, thermal motion significantly perturbs the ion-cavity interaction, leading to enhanced dephasing and a degradation of the EIT interference.